# Search-based DNN Testing and Retraining with GAN-enhanced Simulations

### v2

@@ Abstract
We conducted an empirical investigation with a state-of-the-art, third-party DNN for road segmentation. Our results show that DESIGNATE outperforms TACTIC [33] and DeepJanus [43], two state-of-the-art DNN testing approaches, and several baselines including random search, single objective search, and multi-objective search without GANs. DESIGNATE generates inputs leading to the worst DNN reliability and maximizes input diversity. Further, we demonstrate that the inputs generated by DESIGNATE better support DNN retraining, compared to competing approaches, with an increase in DNN accuracy of eight percentage points. ††margin: R3.6 In addition to our road segmentation case study, in the context of the Mars rover self-driving system, we also demonstrate the applicability and benefits of our approach to object detection on Mars’ surface, using the AI4MARS dataset [50] and a custom-built simulator leveraging Unreal Engine [16] and Airsim. Based on our results across both case studies, we conclude that GANs should be combined with simulators when testing and improving vision-based DNNs as it leads to more effective and diverse testing, as well as better retraining.
@@ II-C2 MarsSim
††margin: R3.6Focusing on extraterrestrial object detection, specifically on the Martian surface, we built a simulator that generate images resembling the ones in the AI4MARS dataset by combining Unreal Engine with the the same plugins that Airsim uses. This simulator creates realistic images of Mars terrain, which, when enhanced with a GAN, enables the generation of diverse training data for testing DNNs in Mars exploration contexts. Including the Airsim API enables us to control the simulator by changing the position of the camera. To create an authentic Martian environment, we modeled objects that match those found in the AI4MARS dataset: rocks, bedrock, sand, and Martian soil, each designed with textures and colors reflective of actual Mars imagery. We also incorporated variations in object size, shape, and placement to reflect the irregular distribution of terrain features typically observed on Mars. Additionally, we introduced terrain undulations and slopes to simulate the topographical challenges of Mars, such as craters, cliffs, and valleys, which pose unique navigational challenges for autonomous rovers.
@@ III The DESIGNATE Approach
GAN-based input generation component. It receives as input the configuration parameter values to be used to generate, with a simulator, scene data (i.e., an image and the corresponding segmentation map). The scene data is then used to generate a realistic image with a GAN trained on a specific reference domain. For our automotive case study, we used the Pix2PixHD GAN trained on the CityScape dataset to generate realistic road scenes from the segmentation map produced by the simulator. ††margin: R3.6 For our Mars case study, we utilized the AI4MARS dataset to train the GAN, and the simulator was built using Unreal Engine and Airsim to generate Martian surface images.
@@ III-B1 Accuracy fitness
The variable carPixelsProportion is computed on the segmentation ground truth. We set the upper bound to to ensure that the car does not cover most of the image, which would result in unrealistic scenarios for evaluating DNN performance. Indeed, images where carPixelsProportion exceeds 0.4 often involve close-ups or unusual viewpoints, such as a vehicle occupying most of the camera’s field of view—lidar-based sensors typically prevent autonomous cars from being in such situations. We set the lower bound above zero to ensure that the image contains at least one car. ††margin: R3.6 Similarly, for the Martian environment, we need to ensure that the sky does not cover most of the image and, therefore, we consider an image relevant if the proportion of pixels belonging to the sky (i.e., ) is within the range:
@@ III-C Search algorithm
Algorithm 2 presents the pseudocode of the NSGA-II algorithm (with archive) employed in DESIGNATE. After creating an empty archive (Line 3), since the initial selection of seeds may affect the search results and inspired by our previous work [3], we generate multiple ( in our experiments) populations of randomly selected individuals (Line 4) and we select the population as the one with the individual having the lowest (Line 6). ††margin: R3.1 This choice is based on empirical observations from prior studies [19], which demonstrated that using multiple seeds helps to reduce bias in the optimization process and ensures a broader exploration of the solution space. For instance, using five seeds strikes a balance between computational efficiency and the reliability of search results in related search-based testing tasks. All the individuals in are then used to update the archive based on Algorithm 1 (Line 7). In practice, the archive is populated with individuals whose distance is above and lead to worst accuracy.
@@ IV-A Subjects of the Study
For the urban environment case, DeeplabV3 was initially trained by its developers on the Cityscapes dataset, which contains a variety of urban street scenes from different European cities. The dataset consists of a variety of images captured at a resolution of 1024 by 2048 pixels in the streets of 50 different European cities, primarily in Germany. Each real-world image comes with a segmentation mask representing its ground truth. The DeeplabV3 training set consists of 2975 images; however, its test set comes without ground truth and, therefore, for testing, we rely on the Cityscapes validation set (500 images). ††margin: R3.3 The average metric (see Section III) is particularly high: 71% (min=00%, first-quartile=62%, third-quartile=91%, max=97%).
@@ IV-B Assessed techniques
In our experiments, we assess a DESIGNATE prototype, that relies on ††margin: R2.1 the original pix2pixHD GAN [38] and AirSim 1.8.1 [37] for the urban case study, and Unreal Engine 5 for the Martian case study. We also extended the NSGA-II algorithm provided by the Pymoo library [7].
To address our RQs, we compare DESIGNATE with a random baseline, two variants of DESIGNATE (for the purpose of performing ablation studies), ††margin: R2.1 TACTIC [33], and DeepJanus [43]. We selected DeepJanus because it may lead to better results if a DNN tends to fail where model behaviour changes. For DeepJanus, in addition to the original implementation provided by its authors, we developed an extended version that integrates a GAN, as described below.
@@
TACTIC is a state-of-the-art DNN testing approach that leverages a MUNIT GAN trained to alter input images by changing their weather conditions (night, sunshine, rain, snow); further, it integrates a search-based approach which alters the latent space vector in such a way that the generated images lead to the worst accuracy for the DNN under test. To apply TACTIC, we randomly select 220 images from the Cityscapes validation set and apply the five weather conditions supported by TACTIC (night, rain, snow night, snow daytime, and sunny conditions), each leading to a distinct new test image. This process generates a total of 1100 images. ††margin: R3.6 Since TACTIC was originally trained to generate driving conditions, we consider it for the Urban environment only; indeed, it makes little sense to add snow and rain effects on Mars images, as well as, changing illumination to make it similar to Earth.
DeepJanus is multi-objective search-based approach that searches for the boundary of the input space that is properly processed by the DNN; specifically, it looks for pair of images that are similar but leads to different DNN results, one of which is erroneous. ††margin: R2.1 We integrate DeepJanus with our simulators with the same configurations as in the original paper. Furher, since DeepJanus exclusively produces simulated images, we explore the impact of using a GAN to generate more realistic images (hereafter, DeepJanusGAN). To achieve this, we incorporated the Pix2pixHD model, as utilized in DESIGNATE, into the DeepJanus workflow to transform simulated images into realistic ones; precisely, DeepJanusGAN, during search, still generates simulator images and ground truth as in DeepJanus, but, instead of providing a simulator image to the DNN under test, it first generates a realistic image using Pix2pixHD and then feeds it to the DNN under test.
@@ IV-C1 Design and measurements
††margin: R3.6For each assessed technique, we compute the , for the urban case study, and for the Martian case study (see Section III-B1 for definitions), obtained by the DeepLabV3 model with the generated images. Since we aim to discover critical DNN limitations, the most effective technique is the one leading to the lowest median, minimum, and first quartile.
@@ IV-C2 Results
Accuracy. ††margin: R3.6 Table II and IV provide, for each of the selected approaches, details on the performance of the subject DNNs for the urban and the Martian environment, respectively. Column images provides the average number of test images selected over the ten executions; for DESIGNATE and its variants we report the images in the archive, for Random and TACTIC we report all the generated images, excluding, for the former, an average of 15 and 169 images, for the urban and Martin environments, respectively, that do not satisfy ‘’. The other columns provide descriptive statistics for and ††margin: R3.6 . Table VII and Table V report differences between each pair of techniques by providing p-values and for the urban and martian environments, respectively. We discuss the results observed with the urban environment first.
@@
Considering pixel-based distance, all the DESIGNATE variants have the same average distance but DESIGNATE still performs better than the others, with a small to high effect size. Such results confirm our observations above for DESIGNATEsingle and DESIGNATENoGAN. Further, the comparison with DESIGNATEpix indicates that a fitness driven by pixel-based distance is less effective than a feature-based distance fitness to maximize pixel-based diversity across images. This indicates that, for the search algorithm, it is easier to generate pixel-wise diverse images when driven by a feature-based fitness. Note that augmenting feature-based distance implies changing the elements depicted in the images and, consequently, the pixel distance. This result may thus be explained by the fact that most of the changes in simulator inputs lead to pixel-based distances that are significant (i.e., above ), whereas only pixel changes that have a strong impact on the elements in the images can result into a significant feature-based distance. Consequently, feature-based distance simplifies the task of the search algorithm to explore the input space towards the search objectives. Also, DESIGNATE performs significantly better than TACTIC and Random, with small and high effect size, respectively. As per feature-based distance, Random presents lower diversity than DESIGNATE (ranges are [0.3,0.76] for Random vs. [0.14,1.00] for DESIGNATE), thus confirming that random sampling is not sufficient to identify the unsafe parts of the input space. TACTIC’s third quartile shows that only 25% of TACTIC’s image pairs have a pixel-based distance higher than 0.68 (i.e., 68% of the pixels differ) while 95% (5th percentile) of the image pairs generated by DESIGNATE present a distance higher than 0.94 (almost all the pixels differ). ††margin: 2.1 Last, the pixel-based distance of DESIGNATE images is significantly larger than that of DeepJanus and DeepJanusGAN images, with medium to large effect sizes, thus confirming the differences discussed above.
@@ IV-D1 Design and measurements
We use the images generated by the approaches considered for RQ1 to retrain our case-study subject DNN. Further, to account for non-determinism, we repeat the process 10 times. For each approach and each retraining iteration, we use the images generated in one run of the approach to augment the Cityscapes training set (2975 images), ††margin: R3.6 for the urban environment, and the AI4MARS training set (16064 images), for the Martian environment. We then retrain the DNN using the augmented dataset, maintaining the same configuration settings (100 epochs) as those used by the DeeplabV3 developers [11]. To avoid introducing bias due to a different number of images in the augmented datasets, for each approach, we randomly select up to 900 images from the set of images generated in RQ1. In addition to the approaches presented in RQ1 (TACTIC, random baseline, DeepJanus, and DeepJanusGAN), we compare DESIGNATE to the training of the DeeplabV3 model for 100 additional epochs using the whole original training set, to determine the degree of improvement achievable with a longer training time, without generating additional training images.
@@ IV-D2 Results
††margin: R3.6Tables X and XII provide descriptive statistics for and from testing the retrained models in the urban and Martian environments, respectively. These tables summarize the performance improvements achieved through retraining with outputs generated by different approaches. Tables XI and XIII present the p-values and values from statistical comparisons of and across retrained models, capturing the significance and effect sizes of the observed differences.
@@
Concluding, our results show that DESIGNATE leads to the best retraining results, thus demonstrating that (1) simulator images are not as effective as realistic images to retrain DNNs, ††margin: R3.6 (2) relying on images belonging to the frontier of DNN behaviours is not the best approach for DNN retraining, and (3) relying on GANs to generate images that simply change weather conditions is not as effective as combining GAN with a simulator to generate realistic images for DNN retraining.
@@ IV-E Threats to validity
To address internal validity threats, we carefully tested our implementation of DESIGNATE and its variants, and further relied on the original version of TACTIC and DeepJanus provided by their authors. ††margin: R3.7 Still related to internal validity, to prevent the generation of unrealistic images with GANs, our approach integrates a strategy to eliminate images that are likely to be unrealistic (i.e., images where the DNN performs better with the simulated image). Such heuristic seems effective since RQ2 results show that retraining the DNN with DESIGNATE images leads to improved results with a real-world test set. If DESIGNATE images were not realistic, the performance of the DNN would have shown limited or no improvement because it would have learned from images that are unrealistic, which is what happens with approaches without GAN. However, it is still possible that our filter may not be fully effective. In the future, we aim to extend our approach with additional methods such as histogram analysis [54] and other state-of-the-art approaches [59, 44] for the identification of out-of-domain and out-of-distribution images.

### v3

@@ Abstract
We conducted an empirical investigation with a state-of-the-art, third-party DNN for road segmentation. Our results show that DESIGNATE outperforms TACTIC [33] and DeepJanus [43], two state-of-the-art DNN testing approaches, and several baselines including random search, single objective search, and multi-objective search without GANs. DESIGNATE generates inputs leading to the worst DNN reliability and maximizes input diversity. Further, we demonstrate that the inputs generated by DESIGNATE better support DNN retraining, compared to competing approaches, with an increase in DNN accuracy of eight percentage points. In addition to our road segmentation case study, in the context of the Mars rover self-driving system, we also demonstrate the applicability and benefits of our approach to object detection on Mars’ surface, using the AI4MARS dataset [50] and a custom-built simulator leveraging Unreal Engine [16] and Airsim. Based on our results across both case studies, we conclude that GANs should be combined with simulators when testing and improving vision-based DNNs as it leads to more effective and diverse testing, as well as better retraining.
@@ II-C2 MarsSim
Focusing on extraterrestrial object detection, specifically on the Martian surface, we built a simulator that generate images resembling the ones in the AI4MARS dataset by combining Unreal Engine with the the same plugins that Airsim uses. This simulator creates realistic images of Mars terrain, which, when enhanced with a GAN, enables the generation of diverse training data for testing DNNs in Mars exploration contexts. Including the Airsim API enables us to control the simulator by changing the position of the camera. To create an authentic Martian environment, we modeled objects that match those found in the AI4MARS dataset: rocks, bedrock, sand, and Martian soil, each designed with textures and colors reflective of actual Mars imagery. We also incorporated variations in object size, shape, and placement to reflect the irregular distribution of terrain features typically observed on Mars. Additionally, we introduced terrain undulations and slopes to simulate the topographical challenges of Mars, such as craters, cliffs, and valleys, which pose unique navigational challenges for autonomous rovers.
@@ III The DESIGNATE Approach
GAN-based input generation component. It receives as input the configuration parameter values to be used to generate, with a simulator, scene data (i.e., an image and the corresponding segmentation map). The scene data is then used to generate a realistic image with a GAN trained on a specific reference domain. For our automotive case study, we used the Pix2PixHD GAN trained on the CityScape dataset to generate realistic road scenes from the segmentation map produced by the simulator. For our Mars case study, we utilized the AI4MARS dataset to train the GAN, and the simulator was built using Unreal Engine and Airsim to generate Martian surface images.
@@ III-B1 Accuracy fitness
The variable carPixelsProportion is computed on the segmentation ground truth. We set the upper bound to to ensure that the car does not cover most of the image, which would result in unrealistic scenarios for evaluating DNN performance. Indeed, images where carPixelsProportion exceeds 0.4 often involve close-ups or unusual viewpoints, such as a vehicle occupying most of the camera’s field of view—lidar-based sensors typically prevent autonomous cars from being in such situations. We set the lower bound above zero to ensure that the image contains at least one car. Similarly, for the Martian environment, we need to ensure that the sky does not cover most of the image and, therefore, we consider an image relevant if the proportion of pixels belonging to the sky (i.e., ) is within the range:
@@ III-C Search algorithm
Algorithm 2 presents the pseudocode of the NSGA-II algorithm (with archive) employed in DESIGNATE. After creating an empty archive (Line 3), since the initial selection of seeds may affect the search results and inspired by our previous work [3], we generate multiple ( in our experiments) populations of randomly selected individuals (Line 4) and we select the population as the one with the individual having the lowest (Line 6). This choice is based on empirical observations from prior studies [19], which demonstrated that using multiple seeds helps to reduce bias in the optimization process and ensures a broader exploration of the solution space. For instance, using five seeds strikes a balance between computational efficiency and the reliability of search results in related search-based testing tasks. All the individuals in are then used to update the archive based on Algorithm 1 (Line 7). In practice, the archive is populated with individuals whose distance is above and lead to worst accuracy.
@@ IV-A Subjects of the Study
For the urban environment case, DeeplabV3 was initially trained by its developers on the Cityscapes dataset, which contains a variety of urban street scenes from different European cities. The dataset consists of a variety of images captured at a resolution of 1024 by 2048 pixels in the streets of 50 different European cities, primarily in Germany. Each real-world image comes with a segmentation mask representing its ground truth. The DeeplabV3 training set consists of 2975 images; however, its test set comes without ground truth and, therefore, for testing, we rely on the Cityscapes validation set (500 images). The average metric (see Section III) is particularly high: 71% (min=00%, first-quartile=62%, third-quartile=91%, max=97%).
@@ IV-B Assessed techniques
In our experiments, we assess a DESIGNATE prototype, that relies on the original pix2pixHD GAN [38] and AirSim 1.8.1 [37] for the urban case study, and Unreal Engine 5 for the Martian case study. We also extended the NSGA-II algorithm provided by the Pymoo library [7].
To address our RQs, we compare DESIGNATE with a random baseline, two variants of DESIGNATE (for the purpose of performing ablation studies), TACTIC [33], and DeepJanus [43]. We selected DeepJanus because it may lead to better results if a DNN tends to fail where model behaviour changes. For DeepJanus, in addition to the original implementation provided by its authors, we developed an extended version that integrates a GAN, as described below.
@@
TACTIC is a state-of-the-art DNN testing approach that leverages a MUNIT GAN trained to alter input images by changing their weather conditions (night, sunshine, rain, snow); further, it integrates a search-based approach which alters the latent space vector in such a way that the generated images lead to the worst accuracy for the DNN under test. To apply TACTIC, we randomly select 220 images from the Cityscapes validation set and apply the five weather conditions supported by TACTIC (night, rain, snow night, snow daytime, and sunny conditions), each leading to a distinct new test image. This process generates a total of 1100 images. Since TACTIC was originally trained to generate driving conditions, we consider it for the Urban environment only; indeed, it makes little sense to add snow and rain effects on Mars images, as well as, changing illumination to make it similar to Earth.
DeepJanus is multi-objective search-based approach that searches for the boundary of the input space that is properly processed by the DNN; specifically, it looks for pair of images that are similar but leads to different DNN results, one of which is erroneous. We integrate DeepJanus with our simulators with the same configurations as in the original paper. Furher, since DeepJanus exclusively produces simulated images, we explore the impact of using a GAN to generate more realistic images (hereafter, DeepJanusGAN). To achieve this, we incorporated the Pix2pixHD model, as utilized in DESIGNATE, into the DeepJanus workflow to transform simulated images into realistic ones; precisely, DeepJanusGAN, during search, still generates simulator images and ground truth as in DeepJanus, but, instead of providing a simulator image to the DNN under test, it first generates a realistic image using Pix2pixHD and then feeds it to the DNN under test.
@@ IV-C1 Design and measurements
For each assessed technique, we compute the , for the urban case study, and for the Martian case study (see Section III-B1 for definitions), obtained by the DeepLabV3 model with the generated images. Since we aim to discover critical DNN limitations, the most effective technique is the one leading to the lowest median, minimum, and first quartile.
@@ IV-C2 Results
Accuracy. Table II and IV provide, for each of the selected approaches, details on the performance of the subject DNNs for the urban and the Martian environment, respectively. Column images provides the average number of test images selected over the ten executions; for DESIGNATE and its variants we report the images in the archive, for Random and TACTIC we report all the generated images, excluding, for the former, an average of 15 and 169 images, for the urban and Martin environments, respectively, that do not satisfy ‘’. The other columns provide descriptive statistics for and . Table VII and Table V report differences between each pair of techniques by providing p-values and for the urban and martian environments, respectively. We discuss the results observed with the urban environment first.
@@
Considering pixel-based distance, all the DESIGNATE variants have the same average distance but DESIGNATE still performs better than the others, with a small to high effect size. Such results confirm our observations above for DESIGNATEsingle and DESIGNATENoGAN. Further, the comparison with DESIGNATEpix indicates that a fitness driven by pixel-based distance is less effective than a feature-based distance fitness to maximize pixel-based diversity across images. This indicates that, for the search algorithm, it is easier to generate pixel-wise diverse images when driven by a feature-based fitness. Note that augmenting feature-based distance implies changing the elements depicted in the images and, consequently, the pixel distance. This result may thus be explained by the fact that most of the changes in simulator inputs lead to pixel-based distances that are significant (i.e., above ), whereas only pixel changes that have a strong impact on the elements in the images can result into a significant feature-based distance. Consequently, feature-based distance simplifies the task of the search algorithm to explore the input space towards the search objectives. Also, DESIGNATE performs significantly better than TACTIC and Random, with small and high effect size, respectively. As per feature-based distance, Random presents lower diversity than DESIGNATE (ranges are [0.3,0.76] for Random vs. [0.14,1.00] for DESIGNATE), thus confirming that random sampling is not sufficient to identify the unsafe parts of the input space. TACTIC’s third quartile shows that only 25% of TACTIC’s image pairs have a pixel-based distance higher than 0.68 (i.e., 68% of the pixels differ) while 95% (5th percentile) of the image pairs generated by DESIGNATE present a distance higher than 0.94 (almost all the pixels differ). Last, the pixel-based distance of DESIGNATE images is significantly larger than that of DeepJanus and DeepJanusGAN images, with medium to large effect sizes, thus confirming the differences discussed above.
@@ IV-D1 Design and measurements
We use the images generated by the approaches considered for RQ1 to retrain our case-study subject DNN. Further, to account for non-determinism, we repeat the process 10 times. For each approach and each retraining iteration, we use the images generated in one run of the approach to augment the Cityscapes training set (2975 images), for the urban environment, and the AI4MARS training set (16064 images), for the Martian environment. We then retrain the DNN using the augmented dataset, maintaining the same configuration settings (100 epochs) as those used by the DeeplabV3 developers [11]. To avoid introducing bias due to a different number of images in the augmented datasets, for each approach, we randomly select up to 900 images from the set of images generated in RQ1. In addition to the approaches presented in RQ1 (TACTIC, random baseline, DeepJanus, and DeepJanusGAN), we compare DESIGNATE to the training of the DeeplabV3 model for 100 additional epochs using the whole original training set, to determine the degree of improvement achievable with a longer training time, without generating additional training images.
@@ IV-D2 Results
Tables X and XII provide descriptive statistics for and from testing the retrained models in the urban and Martian environments, respectively. These tables summarize the performance improvements achieved through retraining with outputs generated by different approaches. Tables XI and XIII present the p-values and values from statistical comparisons of and across retrained models, capturing the significance and effect sizes of the observed differences.
@@
Concluding, our results show that DESIGNATE leads to the best retraining results, thus demonstrating that (1) simulator images are not as effective as realistic images to retrain DNNs, (2) relying on images belonging to the frontier of DNN behaviours is not the best approach for DNN retraining, and (3) relying on GANs to generate images that simply change weather conditions is not as effective as combining GAN with a simulator to generate realistic images for DNN retraining.
@@ IV-E Threats to validity
To address internal validity threats, we carefully tested our implementation of DESIGNATE and its variants, and further relied on the original version of TACTIC and DeepJanus provided by their authors. Still related to internal validity, to prevent the generation of unrealistic images with GANs, our approach integrates a strategy to eliminate images that are likely to be unrealistic (i.e., images where the DNN performs better with the simulated image). Such heuristic seems effective since RQ2 results show that retraining the DNN with DESIGNATE images leads to improved results with a real-world test set. If DESIGNATE images were not realistic, the performance of the DNN would have shown limited or no improvement because it would have learned from images that are unrealistic, which is what happens with approaches without GAN. However, it is still possible that our filter may not be fully effective. In the future, we aim to extend our approach with additional methods such as histogram analysis [54] and other state-of-the-art approaches [59, 44] for the identification of out-of-domain and out-of-distribution images.